The number which best completes the sequence below is:
100 52 28 16 10

Answer by ramkikk66(644)

The number which best completes the sequence below is:
100 52 28 16 10
Ans: 7
Each number is the previous number divided by 2, plus 2.
e.g. 100/2 = 50. 50 + 2 = 52
52 / 2 = 26. 26 + 2 = 28
So next number = (10/2) + 2 = 7
:)